OpenStudy (anonymous):

Solve: 7x − 2(x + 1) = 6x + 14 −12 12 −16 16

OpenStudy (anonymous):

a) multiply out on the left using the distributive property first b) then combine like terms (on the left) c) subtract \(5x\) from both sides, d) subtract 14 from both sides

OpenStudy (anonymous):

that is the method, let me know if you want the answer as well, but try it first

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

off by a minus sign

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

\[7x − 2(x + 1) = 6x + 14 \]multiply out \[7x-2x-2=6x+14\]combine like terms \[5x-2=6x+14\]subtract \(5x\) from both sides \[-2=x+14\]subtract 14 from both sides \[-2-14=x\] \[-16=x\]

OpenStudy (anonymous):

yes, \(-16\)